Q-ACSM 140-146 Chacaltaya station, Bolivia The time in this logbook is writen in Local time Instrument was adjusted to UTC time!! 19/02/2018 The instrument was taken from the box. The stainless steel inlet was bent 45 degrees and the critical orifice was bent too. So, the critical orifice was replaced by a 130 um. Pumps were started. 20/02/2018 Power break for 6 days. 26/02/2018 Pumps were started. 27/02/2018 DAQ was crashing every 10 min. So Phil, from ARI, updated the firaware remotely. Since then, DAQ has not crashed. There was a short between the Filament and the Vaporizer, clearly, the filament was switching off itself everytime the vaporizer was on. So, the instrument was vented and, indeed, the ion cage was touching the vaporizer. Openning a slide screw we pull the vaporizer down and made sure no wires were touching. Pumps were started. 28/02/2018 Flowrate calibration (critical orifice was in) New calibratio values: Y=2.2079 X - 0.58 Room temperature: ~20C The critical orifice was cleaned. Inlet pressure=1.27 Torr 01/03/2018 Lens alignment with AN, 300 nm particles, ~85 #/cm3 OBS: The inlet + lens system looked visually miss-aligned, maybe because of the problem at the customs. New position: Horiz= 8,43 Vertic=8,19 Critical orifice was cleaned. 17:47 - Start measurements with Magee dryer on and flow controller box at 2LPM (1.80 volts). Pressure: Tuning: it improved the shoulder on the right side of n2. Vaporizer Set V=5.7V Heater Bias=89.60V Ions Ref=97 Cathode=-70 Focus=-10.5 Field Axis=-15.0 Extraction=-114 SEM V=1675 V Resolution=65 21:08 - replaced the flow controller box by the water CPC 3787. This model measures down to 5 nm, whilw the CPC 3772 (Tropos) measures down to 10 nm. AB=230 nAmp 22:35 - changed the resolution to 50. 23:08 - the CPC was plugged straight to the main inlet, bypassing the dryer and the ACSM. No change. The water CPC continues measuring 50% less particles. The ACSM was stopped. 23:15 - stop CPC measurements. 23:19 - start ACSM with dryer and flow controller box 23:30 - m/z calibration, new offset=+0.0005 and slope=1.0025 02/03/2018 9:00 Mass calibration. The Calibration system was set up following the diagram from the Italian ACSM. The sillica dryer was replaced by the USP one, using sillica that was dryed the day before, and the nebulizer as well, because the yellow one didn´t have a tube to go into the solution. Since the water CPC was measuring very low concntrations compared to the Total butanol CPC, a butanol CPC was borrowed for the calibration, but the concentrations were still very low for the amount of NH4NO3 added to the solution. Therefore, the nebulizer was once again replaced by the yellow one borrowing the bottle tube from the USP nebulizer and the counts increased immediatly, so the problem might have been the low pressure from the USP nebulizer to the DMA. After that, the calibrations were performed starting from ~900 particles/cc down to zero (water), and the resulting IE amps/loading slope is 2.17e-10. 14:15 Returned to ambient measurements. 07/03/2018 21:39 The SEMV was increased to 1725V. 08/03/2018 10:17 An HEPA filter was connented after the splitter for zero calibration. 18:09 Zero measurements stopped. Restarted ambient measurements. Air pressure dropped to ~0.5 torr on 07/03/2018 around 21:00 UTC but not power cut was registered 09/03/2018 it seems that the critical orifice is dirty. 12/03/2018 14;22 stop measurements and close the valve to clean the criticl orifice, air pressure 0.88torr 14;45 Critical orifice cleaned and air pressure goes to 1.273torr 14;47 zero/filter test start 19;15 filter/zero test end 19;37 ambient measurements start with 0.970tor of air pressure 14/03/2018 fher 14;30 stop measurements to measure the Flow controller flowrate with gilibrator for Vout=1.82V --> Qs=2.05lpm --> Qa=3.99lpm for Vout=2.07V --> Qs=2.66lpm --> Qa=5.10lpm where ´´Qs´´ is flowrate at standard conditions and ´´Qa´´ is florate at ambient conditions 15;38 PM2.5 Cyclone connected 15;50 critical orifice cleaned, current air pressure=1.270torr 16;00 Ambient measurements started with the PM2.5 cyclone setup 2018-03-21 fher (remote) 9:30 system shutdown due to a programed power cut 15:30 System start after the programed power cut. From the begin to the end of the programed power cut a gasoline generator was working at the station. 2018-03-22 fher (remote) 9:00 system shutdown due to a programed power cut 16:30 System start after the programed power cut. From the begin to the end of the programed power cut a gasoline generator was working at the station. 2018-03-23 Fher 14:00 Control box on, then MD1on and When the pressure was 0.09 torr the turbo pumps were turned on 2018-03-24 fher Liine start the measurements but with wrong Voltages 03:18 I have changed the Voltages to the values comming from calibrations. 12:15 HB , SEM tuned /LH 2018-03-26 fher, vale, liine IE calibration performed with 300nm AN. RF = 1.44e-10, RIE(NH4) = 5.74 20:10 Critical orifice cleaned due to the air pressure was 0.56tor, after cleaning is 1.234tor. 20:22 Ambient measrements start 23:28 From 20:22 to 23:20 the number of set was wrong configurated. It was change to 1 sample 1 filter and 28 number of set. 2018-03-27 /LH Critical orifice cleaned, pressure dropped again overnight below 1 torr. After cleaning: p(inlet) = 1.067 torr (inlet open) flow = 1.776 cc/s (inlet open) T(oven) = 636 C (5.7V) SEM = 1771 v (N2 height 1.13e-07) increased after orifice cleaning, now too high. -> SEM tuning -> SEM 1759 V (N2 height 9.67e-08) PUMP stats (inlet open): P1: 650 Hz, 62 W P2: 1500 Hz, 5W P3: 1000 Hz, 5 W fil em: 1 mA m/z cal: 27.974/128.051 17:31 strt ambient measurements having the inlet fully black conductive tubing (had some less appropriate tubing in the line) 28.03.2018 p(inlet) = 1.077 torr flow = 1.788 cc/s 29.03.2018 m/z cal 27.976/128.025 SEM: 1764 V (N2 height 9.50e-08) T(oven) = 638C (5.7V) Pumps stats (inlet open): P1: 650 Hz, 64W P2: 1500 Hz, 5W P3: 1000 Hz, 4-7W (unstable reading) fil em: 1 mA p(inlet) = 1.065 torr flow = 1.761 cc/s 14:11 measurement stopped for flow cal (without critical orifice!), new variables (s = 2.3659, i = -0.58362) saved in DAQ Critical orifice seems to be clogged again.. Need to open and clean. Sulfate calibration performed, RIE(SO4) = 0.83 17:41 UTC back to ambient measurement. 2018-04-02 Fher, Vale, Liine 13:55 Italian ACSM off, therefore the flow was increased to 5 l/min to satisfy the required flow for the PM2.5 cyclone 14:00 Zero test start 16:19 stop zero test 16:20 start ambient measurements 16:34 ACSM hard shut down. Power board it was connected to was shorted. 16:38 Power re-established. ACSM pumping down. Cant figure out how to use the control software. 2018-04-09 Liine 17:38 measurement stopped for IE cal (using right diam particles), critical orifice cleaned. 19:23 back to ambient measurement and new RF(NO3), RIE(NH4) and RIE(SO4) saved in DAQ. RF(nO3) = 9.195e-11 RIE(NH4) = 5.07 RIE(SO4) = 0.67 2018-04-16 Diego, Liine 11:40 critical orifice cleaned. Likely a lot of dust in the air that got it pretty clogged. 11:48 measurement back on. 2018-04-17 Liine 14:20 disconnecting the dryer to see if NO3 and org go up. We are thinking we might lose some semi-volatile stuff in the dryer. I disconnected the dryer power cable too. 14:35 critical orifice cleaned. back to ambient measurement. 2018-04-19 Liine 15:00 critical orifice cleaned, inlet p went up to 1.2 torr, flow 2.25 ccs. PM2.5 cyclone flow checked. It is ok. 16:25 changed vaporizer voltage back to a lower value. I am worried of the refractory stuff (neg spikes in diff spectra). 2018-04-23, 14:24 - Samara Inlet line was unplugged at 14:41 Inlet pressure was 1.05 torr Because the Tee part is not at CHC, we´ll start with the organic standards. m/z calibration: old offset: 0.0005 old slope: 1.0025 old center m/z 28: 27.977 old center m/z 128: 128.042 new offset: 0 new slope: 1.00275 new center m/z 28: 27.992 new center m/z 128: 128.013 15:58 SEMV value increased to 1900 (airbeam too low) 15:59 Malonic acid start ~150 #/cm3 300 nm averaging time: 10 mins, 200 ms/amu -> single scan mode SEm detector was increased to 1900V Vaporizer 5.5V start time 16:56 - end time 17:28 ~55#/cm3 The critical orifice was cleaned and the inlet pressure was back to 1.22 torr 18:26 Xylitol standard starts 300 nm ~ 90#/cm3 inlet pressure 1.20 torr 19:09 xylitol standard ends. At the end the particles went up to 150 #/cm3 19:12 - start sampling ambient air using the Magee Dryer. 24-04-2018 14:18 The calibrations for AN, previous to this were all wrong because the size was different due to altitude, so the size used during calibration was wrong, SEMV was 1830 V by mistake. It was changed back to 1900V and saved again. Measurements were stopped. The critical orifice was replaced by 120um, however, the inlet pressure is not achieving the 1.2 torr anymore, only 1.07 torr. AN Calibratiopn using the Italian system for calibration and a CPC 3772 300 nm 10 LPM Calibration values: NO3 RF=1.80e-10 SO4 RFapp=8.86e-11 NO4 RIE=5.45 So4 RIE=1.10 Inlet P=1.058 Ref N2=1.31e-7 Ref O2=3.15e-8 Rwf Nap=2.01e-11 Ref T= 38.24 AN is atomized inside the system under same conditions, 300 nm, 200mz/amu, 10 min run, (1 closed+1open)x10 ~650 #/cm3 Start at 16:32 at 17:04 the concentration was decreased to ~290#/cm3 at 17:40 the concentration was descreased to 0-1#/cm3 ends at 18:05 AS is atomized inside the system under same conditions, 300 nm, 200mz/amu, 10 min run, (1 closed+1open)x10 ~460 #/cm3 Start at 19:22 - 19:55 Start filter measurements, however, the filter was in the opposite direction. Do not use these data. 25-04-2018 13:42 new filter/zero start 17:37 Zero test stop and ambient measurements start. 27-04-2018 fher From April 26 at 16:00 to April 27 at 14:00 power outgae at the station 18:10 Samara start the system remotly. 03-05-2018 fher 15;41 air beam pressure = 0.945torr a bit clogged, take it out for cleaning 15;42 stop ambient measurements. 15;58 valve closed air beam pressure = 0.028torr 16:00 critical orifice cleaned now air beam pressure = 1.046torr --> ok and start ambient measurements 19;45 power cut ambient measurements stop and 04-05-2018 Power outage from May at 16:00BT to May 4 at 13:10BT (fher) 18:10 the power was back and the system was started. Measurements on (Samara) 05-05-2018 Liine Another sulfate event AB set again to 1.7e-07 instead of 1e-07 (suggested ref), not sure why running SEM with so high volts is preferred?? 10-05-2018 Valeria 11:15 Critical orifice clogged -> cleaned 11.40 Back to ambient measurements. Inlet pressure: 1.051 Torr 22-05-2018 Liine Interlock had kicked on. Fil and SEM off, inlet closed. P was normal. Maybe pump overheating (error2). Measurement restarted. 2018-05-31 Fher 15:19 air beam pressure 1.016tott 15:20 stop measurements for changing the critical orifice 15:34 130um critical orifice setup, air beam pressure 1.135torr. Ambient measurements start 2018-06-?? Samara Power break 2018-06-29 18:13 Instrument starts, SEMV=1950V 2018-07-12 Valeria 14:45 Critical orifice cleand. Inlt P: 1.152 Torr 15:18 Back to ambient measurements 15:20 Inlet pressure droped to 0.97. Cleaning again. Inlet P: 1.149 Torr 15:40 Back to ambient measurements 17:20 Inlet pressure droped to 1.06. Cleaning again. Changed to 120 um orifice. Inlet P: 1.043 Torr 18:00 Back to 130 um orifice. Inlet P: 1.155 17:42 Back to ambient measurements 2018-07-26 DAQ crashed. Peak shape was funny. New tuning and resolution adjust: Resolution was changed from 65 to 35 SetV=6 Bias V=104 Ion Ref=110 Cathode=-70 Focus/-10.5 Field=-15 Extraction=-114 SEMV= 1980V A new AN cal is needed asap! 2018-08-02 Valeria 15:00 approx. false alarm of power cut, therefore measurements were stopped.Power cut re scheduled for Tuesday 07-08-2018 17:20 critical orifice clean. Inlet P 1.144 15:50 orifice clogged again. Claned. Inlet P : 1.144 torr 19:53 Clogged and cleaned... Inlet P: 1.152 torr 2018-08-09 Valeria The critical orifice and all the line were cleaned and the ACSM is back to work. 2018-08-13 Isabel sulfate event seen in both the ACSM and SO2 (>100 ppb) analyzer. 2018-08-16 Laura Critical orifice got clogged on 14-08-2018 early morning Orifice clened. Inlet P: 1.150 2018-08-23 Isabel Power outage from 18-08-18 14h47 until 21-08-21 17h35. ACSM PC remained off. ACSM turned off as instructed by Marcos Andrade 2018-08-30 17:50 Samara SEMV was reduced to 1880V because it was too high AB. Resolution was decreased to 45. 2018-09-09 Samara Servo is broken Increased the SEM V to 1920V 2018-09-26 Valeria, Junior 17:43 Servo successfully replaced. Inlet valve fixed. Critical orifice cleaned. Inlet P: 1.168 torr It looks lke the problem was related to a blown fuse on the board. It was replaced to 1.5A. M/z cal adjust, new slope=1.0053 and offset=0 2018-10-08 Laura,Fer Air pressure is 1.156 torr --> ok mass data is cakculating in a good way (sample - zero) Instrumento funcionando regularmente. La bomba principal fue apagada entre las 16h07 y 19h15 UTC. 2018-10=15 fher, samarautc power outage from october 11 around 18UTC to october 12 around 18 When power was back ACSM started but PC did not (I do not know why) On October 15 around 14UTC PC was turned on and Samara restarted the adquisition software. Pressure line = 1.168 torr --> ok Around 18UTC I have change: Configuration/digital line configuratio/filter Valve to direct drive 2018-10-16 Fher (remote) ACSM working fine, so much sulfate these days 2018-10-18 Fher 14:30UTC start zero test 18:55 zero test ended the dryer is working fine!!!!! the issue was related to the power of the vaccum pump. The one that Valeria setup is not strong enough to generated an adequate vaccump pressure in the dryer. So, I have changed the pump to a stronger one then I adjusted the vaccum meter to -0.25bar which produce a counterflow of 4.9 lpm which is in a good compromise between sample flow-counterflow-vaccum pressure. air pressure = 1.171 torr --> great ACSM is working fine 2018-10-23 Fher (remote) Power outage from 19 to 20UTC, after that system restarted and start measurements, 2018-10-25 fher 18:14 ACSM ok 2018-10-29 fher, laura 14:15 ACSM ok air pressure = 1.164 torr 2018-11-05 Fher (remote) ACSM is working fine and saving data air beam pressure = 1,167 torr 2018-11-08 Manuel, Isabel 16:19 Instrument is working fine and saving data correctly. air pressure=1,169 torr --> ok (fher) 2018-11=12 fher power outage on nov 10 from 17 to 21UTC ==> no data until today 14:30 Prisma and PC restarted due to a lack of comunication Ambient measurements start at 14:53 SEM voltage chango from 2020 to 1970 System is working and saving data 2018-11-14 fher (remote) 13:50 System fine and saving data 2018-11-15 fher (remote) 13:30 System fine and saving data air beam pressure = 1,166torr --> good 2018-11-16 Fher (remote) UPS failed on Nov 15 and measurements stoped, at 21:10 Measurements restarted, Air beam pressure = 1,158torr --> ok 2018-12-27 Laura, Marcelo Instrument is working ok. 2018-11-19 fher laura 18:53 System fine air pressure 1.163torr 2018-11-21 Fher (remote) 15:43 air pressure = 1,161torr -->ok 15:41 Ambient measurements start 15:34 System restarted after the power outage Power outage on nov 20 from 3 to 21UTC 19:24 after few crashes (DAQ) it is ok. The AB in much higher since Nov 11th. So, the SEMV was decreased to 1900V. 2018-11-26 fher, laura DAQ problem from Nov 24 (afternoon) to nov 26 (morining) 14:30 PC and prisma restarted 14:50 ambient measurements start 16:48 air preassure = 1.174 torr 2018-11-29 Manuel 14:00 instrument is working fine and saving data correctly. Inlet pressure 1.174 torr 2018-12-17 Manuel 16:59 instrument is working fine and saving data correctly. Inlet pressure 1.175 torr 2018-12-24 Isabel Instrument seems fine.1.175 torr 2019-01-10 Manuel 15:55 Instrument is working fine and saving data. Inlet pressure 1.171 torr 2019-01-14 Manuel 15:37 Instrument is working fine and saving data. Inlet pressure 1.176 torr 2019-01-24 fher from De 1 to today there is complete data, critical orifice clean enough (1;172torr) and sample RH lower than 20% 2019-01-28 Samara, fher There was a power break on the 25th. The instruments was restarted. SEM V was decreased to 1900V. 2019-02-01 fher (remote) 14:33 ACSM working fine and saving data. Air beam pressure ok (1.171 torr) 2019-02-04 Laura 14:37 Instrument is working ok. Inlet pressure -> 1.172 torr. 2019-02-07 fher 17:02 ACSM working fine and saving data. Air beam pressure ok (1.170 torr) 2019-02-11 Manuel, Isabel 15:26 instrument is working fine and saving data correctly. Inlet pressure 1.170 torr 2019-02-14 fher, manu System restarted today at 15:00 Power outage on feb 13 from 9 to 18 UTC. 2019-02-18 fher (remote) 20:45 Although it was changed from Gear to Direct drive, the measurements are still negative. 2019-02-25 fher 18:00 ambient measurements start and now w have good data. system shutdown at 17:30 18:05 system started and ambient measurements. It seems that no sample air was measure from last power outage because in scan parameters scan number was set to zero. This time scan and filter were setup to 1 each other. 2019-02-28 fher ACSM working fine air beam pressure = 1.171 torr --> fine 2019-03-06 fher (remote) System working fine 2019-03-07 Fher (remote) 13:07 System working fine, saving data and air beam pressure ok (1.167torr) 2019-03-11 fher 14:20 ACSM working fine and saving data air beam pressure = 1.168 torr --> fine 2019-03-14 fher power outage from march 12 18UTC to March 13 18UTC Laptop did not start when the power back 14:22 ACSM starts, air beam pressure closed valve = 0.041torr 14:30 ambient measurements start 14:32 ACSM crashed --> prisma restarted Laptop time 5 minutes delay from actual time --> corrected at 14:35 (old time) 14:41 ACSM starts 17:59 ACSM working well 2019-03-18 fher (remote) System working fine 2019-03-21 fher (remote) 13:06 System working fine 2019-03-25 fher 17:11 System working fine (1.162torr air pressure), however AB is decreasing so fast (according to Samara) 2019-03-28 fher 14:49 ACSM working fine and saving data air beam pressure = 1.172 torr --> fine 2019-04-08 fher 18:09 ACSM working fine and saving data air beam pressure = 1.162 torr --> fine 2019-04-11 fher on april 10 at 17:36 something triggred the internal protection, could not be a porwer outage so could be an internal issue. 14:40 waring window close and filament, detector turned on and off --> it works and the system start to measure again. Let see next days if the internal protection is activated. air beam pressure 1.167 torr --> ok 2019-04-22 fher 14:33 System ok and saving data air beam pressure 1.160 torr --> ok 2019-05-09 fher 17:33 System ok and saving data air beam pressure 1.164 torr --> ok 2019-05-19 fher 13:37 It seems that the Wifi antenna is gone --> so an external Wifi antenna is connected and now there is internet 13:40 system working fine but air beam pressure is low 1.142 torr 2019-05-24 samara 18:52 The SEMV was increased to 1930V. 2019-05-27 fher 14:40 ACSM conflict with prisma IP from May 24 --> Pc restarted and prisma disconnected and connecte again (after 2 minutes) 14:50 Pc and ACSM software restarted 15:00 ambient measurements 2019-05-30 fher 16:46 air beam pressure = 0.496 torr --> to low 2019-06-03 fher 14:53 air beam pressure = 0.425 torr --> to low 14:55 stop ambient measurements 15:30 critical orifice cleaned, air beam pressure = 1.155torr --> great 15:35 back to amient measurements utc17:31 systm working fine 2019-06-10 fher 17:30 System ok and saving data air beam pressure 1.149 torr --> ok 2019-06-13 Laura,Fer 18:22 Instrument is working ok. inlet pressure: 1.157 , which is ok. 2019-06-24 fher From June 20 at 11:13 to date there was a filament error and no data were took the error was solved at 15:30 and from that ambient measurements is taking air been pressure 1.143torr --> ok 2019-06-27 fher 19:16 ACSM working fine inlet pressure: 1.144 --> ok. 2019-07-01 fher 19:34 ACSM working fine inlet pressure: 1.159 --> ok. 2019-07-04 Isabel 14:33 instrument working fine. TV relaunched. 2019-07-08 fher 14:39 from July 5 at 17UTC main pump crashed, however ACSM measuremenents seems to be fine. ACSM is working fine inlet pressure: 1.154 --> ok. 2019-07-11 fher 17:40 ACSM working fine air beam pressure: 1.149 --> ok. 2019-07-19 Isabel, Laura Erupcion volcan Ubinas 7 utc 2019/07/23 Laura, Isa Instrument working ok. 2019-07-29 fher 15:01 ACSM working fine air beam pressure: 1.154 --> ok. 2019-08-01 Fher 15:49 Instrument is working ok. inlet pressure: 1.153, ok 20:20 m/z 28 less than 100mA, so SEM voltag changed from 1930V to 2000V 2019-08-15 fher 14:00 On Aug 8 around 20UTC an RF temperature error show up and stopped the system. This sissue is related to mass spectrometer temperature. 14:39 prisma filter changed to a new one, the old one was full of dust and it was not possible to clean it. However, the system still shows the RF temperature error. 16:00 It was observed that the internal prisma fun is not running, this could be the reason to have this high temp error. We asked instruction to Manufacturer about how to change this fun 2019-08-22 fher 15:55 Prisma internal fun changed to a new one. In the local market only was possible to find a fun of 12V-0.13A with the same size as the old one. The old one fun is 24V-0.21A; As the new fun needs less voltage and current, it runs without problems. 2019-08-26 fher 18:46 System running well 2019-08-29 fher 18:54 System running well. air beam pressure 1.155torr --> ok 2019-09-05 fher 19:01 System running well. air beam pressure 1.162torr --> ok 2019-09-09 fher 15:00 System running well. air beam pressure 1.159torr --> ok 2019-09-12 Laura, Isa 16:02 Instrument is working ok. Inlet pressure: 1.171 torr --> ok 2019-09-16 fher 14:10 Instrument is working ok. Inlet pressure: 1.145 torr --> ok 15:02 ACSM turned off due to a general maintanace at Chacaltaya. 2020-01-07 fher Power back to Chacaltaya 19:00 controlbox on --> MD1 on --> turbo pumps on --> keep the system producing vacuum until Jan 9 2020-01-09 fher Turbo pumps = 0.07torr 14:10 Prisma connected 15:18 ACSM DAQ has stoped ==> prisma restarted by software 15:22 Software starts and prisma connected 15:25 mz28=750 too high (SEM=2000) --> mz28=430 (SEM=1920) --> mz28=350 (SEM=1880) --> mz28=170 (SEM=1780) --> mz28=115 (SEM=1720) --> ok 17:02 airbeam pressure = 1.171torr ACSM working fine 2020-01-16 fher 14:23 System working fine, low concentration of all variables from Jan 12 probably we are in clouds and wind comming from East air beam pressure 1.165torr --> ok 2020-01-23 fher 19:15 air beam pressure 1.161torr --> ok system ok 2020-01-27 Laura 16:14 Instrument working properly. Air beam pressure 1.162torr --> ok system ok 2020-01-30 Laura, Fernando 19:42 Instrument working properly. Air beam pressure 1.160torr --> ok system ok 2020-02-04 Samara e Lorrany Aumento do SEM-V de 1780 para 1820 15:04 Filter start 15:10 Changing filter time to 10 minutes 15:35 - m/z cal, old values = offset=0, slope=1.0053 New m/z cals offset=-0.0006, offset=1.004 19:22 - start filter measurements at same ambient conditions. There was a ower break just after we left, ignore these filter period 2020-02-06 Samara and Lorrany AN Cal1, 300 nm, using home made calibration system at 6LPM and Buthanol CPC at 1LPM. NO3 RF=3.22e-10 nh4 RIE=5.76 Illet p=1.186 Ref n2=1.47e-7 ref o2=7.12e-10 ref nap= 1.42e-11 ref T= 36.28 AN Cal2 NO3 RF=3.20e-10 nh4 RIE=5.21 Illet p=1.186 Ref n2=1.45e-7 ref o2=6.63e-10 ref nap= 1.45e-11 ref T= 36.60 17:49 - atomizing AS at 300nm varying the concwntrations. 18:45 - stop. It's possible the line was leaking in the first points, not later. 19:26 - start HEPA filter measuerement - zero - using silica gel dryer and same cycle as ambient measurements, using duty cycle: (1+1)x28 ~30 min. 2020-02-07 Samara 13:13 - stop HEPA filter. 13:14 - start measuring room air in 4min cycle.